How will you solve

What is the change in temperature from 127F to 120rCin kalvin?
Answer #1

Here ,for changing 12degree F to C

T(C)/100 = (12 - 32)/180

T(C) = -11.1 degree C

change in temperature in degree C = 120 - (-11.1)

change in temperature in degree C = 131.1 degree C

change in temperature in degree C = change in temperature in K

change in temperature in K = 131.1 K

the change in temperature is 131.1 K